Aliens repair

The first time I put this board onto the workbench to have a look at it, there was a bit of corrosion on a few ICs. So before starting I gave the board a wash with SafeWash and then rinsed it off with Electronic Board Cleaner dried if off thoroughly. This removed the bulk of the corrosion and cleaned all the dirt and grime from the board.

Next I moved onto removing the ICs which had corrosion on them and replaced them. This included replacing G12, G13, G19 and F10.

With the corrosion taken care of, I moved onto the eproms and checked them in the programmer. The IC at E24 tested fine, but the IC in C24 was not even from Aliens. So I pulled out a new 27C010 and programmed a new eprom and installed it.

Powering up the system resulted in a constant reset loop. I tested the voltage at some of the ICs and found that I was only getting about 4.7 volts. The Jamma connector on the board was very badly worn, so I increased the voltage on the bench power supply until I had close to 5v at the logic ICs.

Increasing the voltage stopped the board from resetting so rapidly, which was being caused by the power watchdog built into the 051550 module. However it was still rebooting. So it was time to start at the beginning and ensure that the system clock was running.

Firing up the CRO I was happy to see a steady 24MHz at pin 6 and 8 of IC G11. Next I checked around the 74245 at G19 to see if there was data between the CPU, SRAM and ROMs. Both sides of the 245 appeared to be quite active, so it was time to move onto the SRAM IC at locations G21 and G23 respectively.

The SRAM at G21 appeared to be working fine, but checking around G23 with the logic probe showed no activity at all on any of the data lines. I piggybacked G23 with a new SRAM and probed the data lines again, and this time there was plenty of activity. So I removed the faulty IC and installed the replacement.

After powering up the board again I was greeted with the garbled screen with colour. It appeared as though the board was trying to run, but the video was bad. So I checked around the video SRAM at E14 and E15 and found the ram at E15 to have a number of dead data lines. I replaced the 16kb SRAM and fired up the board once again. This time the board came up with the test screen showing all the RAM and ROMs as being OK. I turned dip switch 3 on bank 3 to off and power cycled the board again and this time I was greeted with the game.

Letting the game run a bit and turning on sound in attract mode I noticed that there appeared to be some sounds missing. So I programmed a new IC for position D5 and replaced the original mask rom and the final issue was resolved.